OpenStudy (anonymous):

ax + 4b > 9 x > the quantity negative 3 times b minus 5 all over negative a x < the quantity 5 minus 3 times b all over negative a x < the quantity negative 3 times b plus 5 all over negative a x > the quantity negative 3 times b plus 5 all over negative a

OpenStudy (anonymous):

@Abhisar

OpenStudy (abhisar):

Hey there! Welcome to OpenStudy :)

OpenStudy (abhisar):

@TheSmartOne

OpenStudy (anonymous):

Thank you @Abhisar

OpenStudy (anonymous):

@Abhisar do you mind helping me still. I am still waiting here for help.

OpenStudy (abhisar):

Sorry Mathematics is not my forte :(

OpenStudy (anonymous):

Oh may you tag someone to help me @Abhisar

OpenStudy (abhisar):

@Jhannybean may help you

OpenStudy (jhannybean):

ax + 4b > 9 Since you're solving for the quantity x, you want to move everything else on to the other side of the \(\gt\) sign

OpenStudy (jhannybean):

So you'll subtract -4b from both sides of the inequality.

OpenStudy (jhannybean):

DO that and tell me what you get.

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

-ax>9 - 4b?

OpenStudy (jhannybean):

Tell me why you made it negative.

OpenStudy (anonymous):

made what negative

OpenStudy (jhannybean):

Why have you written -ax on the left hand side? you're subtracting -4b from the +4b that is already there. No need to change ax

OpenStudy (anonymous):

oh sorry

OpenStudy (jhannybean):

ax + 4b > 9 - 4b -4b ______________________ ax > 9 - 4b This is what you got, yes?

OpenStudy (anonymous):

no wait there is supposed to be a negative in front of ax so it would be -ax I wrote the equation wrong

OpenStudy (jhannybean):

So your original equation is -ax + 4b > 9 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry for the confusion my bad

OpenStudy (jhannybean):

No problem :)

OpenStudy (jhannybean):

So yes, your previous post would be correct.

OpenStudy (anonymous):

ok so now we have to continue

OpenStudy (jhannybean):

-ax > 9-4b \(\checkmark\)

OpenStudy (jhannybean):

When you divide or multiply both sides of an inequality by a negative value, you always have to flip the sign. This means \(\gt\) will become \(\lt\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what will I have to do next?

OpenStudy (anonymous):

@Jhannybean

OpenStudy (jhannybean):

So divide both sides of your inequality by -a, this will flip your inequality. What do you get?

OpenStudy (anonymous):

i REALLY DO NOT KNOW i AM CONFUSED

OpenStudy (jhannybean):

It's ok, it's ok. calm down :)

OpenStudy (anonymous):

oh sorry did not mean to use caps I am trying to type without having my eyes on the keyboard

OpenStudy (jhannybean):

\[\frac{-ax}{-a} \gt \frac{9-4b}{-a}\]Like I mentioned, When dividing by or multiplying by a negative number, in this case -a, the sign will FLIP. so what do you end up with?

OpenStudy (anonymous):

I do not know

OpenStudy (jhannybean):

We can eliminate answer choices A and D.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

this is just a guess is the answer a

OpenStudy (anonymous):

@Jhannybean

OpenStudy (jhannybean):

... didn't we just say the answer choice A and D can be eliminated?

OpenStudy (jhannybean):

eliminated means removed, as in they do not count. As in they are not going to be one of your answer choices,

OpenStudy (anonymous):

oh sorry I am a hot mess today

OpenStudy (anonymous):

I am going to guess on b then

OpenStudy (anonymous):

@JhannyBean

OpenStudy (jhannybean):

Just letting know now, all your answer choices are incorrect for this problem. None of them match up to your function.

OpenStudy (anonymous):

they are let me see

OpenStudy (anonymous):

Solve for x. -ax + 4b > 9 x > the quantity negative 3 times b minus 5 all over negative a x < the quantity 5 minus 3 times b all over negative a x < the quantity negative 3 times b plus 5 all over negative a x > the quantity negative 3 times b plus 5 all over negative a

OpenStudy (anonymous):

is that better

OpenStudy (jhannybean):

\[\frac{-ax}{-a} \gt \frac{9-4b}{-a}\]\[x \lt \frac{9-4b}{-a}\] Or otherwise written as "x < the quantity of 9 minus 4 times b all over negative a"

OpenStudy (jhannybean):

So I guess b would work here, if that's all you're left to work with :\

OpenStudy (anonymous):

may u help me with two more

OpenStudy (jhannybean):

Sorry, I'll tag in someone else for now. Open up another question and post your problem! More people can help :D

OpenStudy (anonymous):

ok